Astrid1:

I don't know what to do, she told she would tell us next week about our assignment, but it's due today, anyways, I need help. Idk if I just don't remember or whatever, anyways I need help.

Astrid1:

Shadow:

One way you could do this is find the area of the black square then subtract from it the area of the triangle (between the black square and the blue squad) times 4.

Astrid1:

Okay... I'm not sure if I'm doing this right, but would that make it 400?

Shadow:

That sounds a bit too large.

Shadow:

The length of the sides of the square are 9 squares long, so you'd calculate the total area of the black square to be 9 times 9 = 81

Astrid1:

YEah.. I just looked at it again. So I would do..\[9\times9\times4\] \[9 \times 9~obviousily = 81.~\And~81\times4=324\] So 324 would be the correct answer?

Shadow:

That would be four black squares.

Shadow:

What were trying to do is find the area of the black square, which is 81, then subtract from that value, the area of the 4 triangles.

Shadow:

This will give us the area of the blue shaded square.

Shadow:

1 attachment
Shadow:

Conceptually this is what we're doing. Black square - the red regions (triangles) = area of the blue square

Shadow:

@Astrid1 Does that make sense?

Astrid1:

Yeah, kind of. I'm still trying to figure out how to find the area of the triangles.

Shadow:

\[A_{t} = \frac{ 1 }{ 2 }bh\] Area of a triangle is equal to 1/2 of the base multiplied by the height

Shadow:

The base is 1 and the height is 9.

Astrid1:

OKay so my answer would be 4.5? Ngl, I used google xd.

Shadow:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow \[A_{t} = \frac{ 1 }{ 2 }bh\] Area of a triangle is equal to 1/2 of the base multiplied by the height \(\color{#0cbb34}{\text{End of Quote}}\) \[A_{t} = \frac{ 1 }{ 2 }(1)(9) = \frac{ 1 }{ 2 } (9) = 4.5\] It's not that bad :P Now since we have the area of one triangle, we multiply that by four.

Astrid1:

So it would be 18?

Shadow:

Yes. So now we have the area of the four triangles. What do we do next?

Astrid1:

I don't know..😅😕

Shadow:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow What were trying to do is find the area of the black square, which is 81, then subtract from that value, the area of the 4 triangles.

Shadow:

Remember the square. We can find the area of the smaller square if we subtract the area of the triangles from the bigger square.

1 attachment
Astrid1:

So 81 minus 18? o-o

Shadow:

Yes

Shadow:

Big square minus triangles equals little square

Astrid1:

63?

Shadow:

Yeah you got this :p